Solve
2^(2x - 1) = 8^(x + 7).

Answer by chibisan(131)

2^3 = 8
2^(2x - 1) = 2^3(x + 7)
2^(2x - 1) = 2^(3x + 21)
since the base is same , the index will be equal
2x - 1 = 3x + 21
-1 - 21 = 3x - 2x
-22 = x